OpenStudy (anonymous):

Please solve the equations by using Substitution and get a free medal! 4x-y=15 -2x+3y=12

OpenStudy (anonymous):

solve for y in the first equation y=4x-15 substitute this for y in the second equation -2x+3y=12 becomes -2x+3(4x-15)=12 -2x+12x-45=12 10x=57 x=57/10

jhonyy9 (jhonyy9):

this is right and y=4(57/10) -15 = 114/5 -15 =(114-75)/5 = 41/5